For the experiment of drawing a single card from a standard 52-card deck, find the probability that the card is neither a heart nor a 7
Answer by ewatrrr(24785)

single card from a standard 52-card deck drawn:
P(neither a heart nor a 7) = 1 - P(H 0r 7) = 1- 16/52 = 36/52 = 9/13